OpenStudy (anonymous):

If (z o z)(x) =1/16x, what is z(x)?

OpenStudy (amistre64):

hmm, assume z = ax for simplicity z(z) = a(ax) = a^2x let a^2x = 1/16x and solve for a maybe?